OpenStudy (anonymous):

Hours Studying and Test Grades Time (hours) Grade 1 66 2.5 70 3 81 3.5 92 4 97 Jennifer believes that she can predict her test scores based on how long she studies. She asks 5 of her friends how long they studied and what grade they made on the test and then preforms a regression to model the data. Her data is shown in the table. Which function best models the data?

OpenStudy (anonymous):

y = 59.576(1.108)x B) y = 63.461(1.105)x C) y = 2.833x2 - 3.833x + 67 D) y = 7.733x2 - 24.4x + 82.667

OpenStudy (anonymous):

Have worked all of them for all points, but the only one that comes close to 97 after 4 hrs of study is C. A and B look odd.